# footnotetext: MSC2020: Primary 34A08

### v1

@@ 1. Introduction
Duhamel’s principle is a classical tool in the theory of linear evolution equations, providing a method to reduce inhomogeneous problems to homogeneous ones. Fourier’s early work on the heat equation laid the analytical foundations for superposition methods and integral representations of solutions. Building on these ideas, Duhamel formulated in the 1830s a general principle for reducing the Cauchy problem for an inhomogeneous linear partial differential equation to the corresponding homogeneous problem. This reduction principle, now known as Duhamel’s principle, was subsequently employed by Kirchhoff in the study of wave propagation, leading to explicit solution formulas for the wave equation. Later, Hadamard placed these representations within a rigorous analytical framework by developing the modern theory of the Cauchy problem for hyperbolic equations, clarifying issues of well-posedness, causality, and the validity of Kirchhoff type formulas. Since then, Duhamel’s principle has become a standard tool in the analysis of linear evolution equations, with applications ranging from parabolic and hyperbolic partial differential equations to dispersive and quantum-mechanical models.
In recent decades, the development of fractional evolution equations has motivated attempts to generalize Duhamel’s principle to nonlocal-in-time operators, such as Caputo and Riemann–Liouville derivatives.
@@
the classical semigroup approach fails due to memory effects: forcing at time $s$ affects all future times nonlocally. Consequently, the classical Duhamel reduction cannot be applied directly. The fractional analogue of Duhamel’s principle was discussed in [25, 26], leading to the following formula:
@@
with the ”fractional impulse” applied to the forcing. Here, each integrand can be viewed as a homogeneous fractional problem “initiated” at time $\tau$ by a fractional impulse, but the evolution now depends on the entire past history rather than only the initial value at $\tau$. The key features of this approach are:
@@
There is a substantial literature on fractional order evolution differential-operator equations. Below, we mention only works that are closely related to the representation of solutions to such equations. Kochubei [15] studied existence and uniqueness of a solution to the abstract Cauchy problem $D_{\ast}^{\alpha}u(t)=Au(t),\,u(0)=u_{0},$ with Caputo-Djrbashian fractional derivative for $0<\alpha<1$ and a closed operator $A$ with a dense domain ${\mathcal{D}}(A)$ in a Banach space. El-Sayed [9] and Bazhlekova [4] investigated the Cauchy problem for $0<\alpha<2.$ In the more general case of $\alpha>0,$ Gorenflo et. al. [12] studied existence of solutions in Roumieu-Beurling and Gevrey classes. Kostin [17] proved that the abstract initial value problem (Cauchy type problem) $D_{+}^{\alpha}u(t)=Au(t),\,D_{+}^{\alpha-k}u(0)=\varphi_{k},\,k=1,...,m,$ where $\alpha\in(m-1,m)$ and $D_{+}^{\alpha}$ is the Rieman-Liouville derivative, is well-posed. For more information about recent results on the Cauchy problem for abstract fractional differential-operator equations, we refer the reader to [4, 8, 14]; and for a recent mathematical treatment of the distributed fractional order differential equations to papers [16, 19, 24].
These approaches to fractional evolution equations, including Bazhlekova [4], Mainardi et al. [3], and Kochubei [15], provide rigorous solution representations and fundamental solutions. However, these methods do not produce a true Duhamel-type reduction: the inhomogeneous solution is not expressed as a superposition of homogeneous problems initiated by fractional impulses. Hence, the “Duhamel philosophy” of reducibility is not realized in these approaches.
The true fractional analog of Duhamel’s principle, in the particular case of fractional order partial differential equations with a single ”time-fractional” term in the evolution equation, was obtained in [25, 26], for distributed order evolution equations in [27]. See also [28, 29] for the survey of the topic and [1, 2, 6, 7, 11, 13, 18, 20, 21, 22, 29, 31] for various applications of the fractional Duhamel principle.
As regards evolution equations involving the fractional derivative in the sense of Riemann–Liouville with classical initial (Cauchy) conditions, such problems are ill-posed. Consequently, the notion of an “initial impulse,” discussed above, is not applicable. Nevertheless, one key feature of the Duhamel principle can be retained. Specifically, in an appropriate formulation, the same solution operator associated with the homogeneous problem can also be used to solve the nonhomogeneous problem.
@@
The operator $A\circ\mathfrak{D}^{\mathscr{A}}$ on the left-hand side of system (2) is defined as the Hadamard (entry-wise) product of the matrix $A=(a_{j,k})_{j,k=1}^{m}$ and the matrix-valued operator $\mathfrak{D}^{\mathscr{A}}=(D^{\alpha_{j,k}})_{j,k=1}^{m},$ where $D^{\alpha_{j,k}}$ denotes the Riemann–Liouville or Caputo–Djrbashian fractional derivative of order $\alpha_{j,k}.$ That is, the left hand side of (2) reads:
@@
The precise conditions on the matrix $A,$ the matrix-valued operator $F,$ and the vector-valued functions $H(t)$ and $\Phi$ that ensure the existence and uniqueness of a solution to problems (2) and (3) will be specified in Section 2, along with a review of the classical Duhamel principle and the basic reference vector spaces used in this paper. In Section 3 we state the main results: an abstract fractional analog of Duhamel’s principle for fully coupled systems, and discuss some of its applications. Since the system is fully coupled, the associated fractional impulses are expected to be coupled as well, and their form of coupling is also addressed.
@@ 3. The fractional Duhamel principle for fully coupled systems
If the system coefficients satisfy $a_{ij}=0$ for $i\neq j,$ the system becomes decoupled in time, and the memory of each component evolves independently of the others. In this case, $Q(t)\equiv 0$ and the fractional impulses depend only on the history of the corresponding components.
Corollary 12.
@@
Conclusion similar to Corollary 12 is valid in the case of systems given by the fractional derivatives in the sense of Riemann-Liouville. In the special case $m=1,$ Theorems 9 and 10 recover the fractional Duhamel principle established in [25, 26].
@@
where $H\in AC^{(}[0,T],L_{1}(\mathbb{R}))^{2}$ and $\Phi(x)\in L_{1}(\mathbb{R})\times L_{1}(\mathbb{R}).$ In equation (31) the time-fractional orders are $\alpha_{11}=1/2,\alpha_{22}=1/3$ and coefficients $a_{12}=a_{21}=0,$ implying $g_{1}(t)=g_{2}(t)=0.$ Hence, in accordance with Corollary 12 the solution of the Cauchy problem (31)-(32) has the representation

### v2

# arXiv:2602.10379v2[math.CA] 28 Feb 2026 footnotetext: MSC2020: Primary 34A08

@@ 1. Introduction
Duhamel’s principle is a classical tool in the theory of linear evolution equations, providing a method to reduce inhomogeneous problems to homogeneous ones. Fourier’s early work on the heat equation laid the analytical foundations for superposition methods and integral representations of solutions. Building on these ideas, Duhamel formulated in the 1830s a general principle for reducing the Cauchy problem for an inhomogeneous linear partial differential equation to the corresponding homogeneous problem. This reduction principle, now known as Duhamel’s principle, was subsequently employed by Kirchhoff in the study of wave propagation, leading to explicit solution formulas for the wave equation. Later, Hadamard placed these representations within a rigorous analytical framework by developing the modern theory of the Cauchy problem for hyperbolic equations, clarifying issues of well-posedness, causality, and the validity of Kirchhoff type formulas. Since then, Duhamel’s principle has become a standard tool in the analysis of linear evolution equations, with applications ranging from parabolic and hyperbolic partial differential equations to nonlinear dispersive and quantum-mechanical models [Sogge, Tao].
In recent decades, the development of fractional evolution equations has motivated attempts to generalize Duhamel’s principle to nonlocal-in-time operators, such as Caputo and Riemann–Liouville derivatives [25, 26, 27].
@@
the classical semigroup approach fails due to memory effects: forcing at time $\tau$ affects all future times nonlocally. Consequently, the classical Duhamel reduction cannot be applied directly. The fractional analogue of Duhamel’s principle was discussed in [25, 26], leading to the following formula:
@@
with the ”fractional impulse” applied to the forcing. Here, each integrand can be viewed as a homogeneous fractional problem initiated at time $\tau$ by a fractional impulse, but the evolution now depends on the entire past history rather than only the initial value at $\tau$. The key features of this approach are:
@@
There is a substantial literature on fractional order evolution differential-operator equations. Below, we mention only works that are closely related to the representation of solutions to such equations. Kochubei [14] studied existence and uniqueness of a solution to the abstract Cauchy problem $D_{\ast}^{\alpha}u(t)=Au(t),\,u(0)=u_{0},$ with Caputo-Djrbashian fractional derivative for $0<\alpha<1$ and a closed operator $A$ with a dense domain ${\mathcal{D}}(A)$ in a Banach space. El-Sayed [8] and Bazhlekova [3] investigated the Cauchy problem for $0<\alpha<2.$ In the more general case of $\alpha>0,$ Gorenflo et. al. [11] studied existence of solutions in Roumieu-Beurling and Gevrey classes. Kostin [16] proved that the abstract initial value problem (Cauchy type problem) $D_{+}^{\alpha}u(t)=Au(t),\,D_{+}^{\alpha-k}u(0)=\varphi_{k},\,k=1,...,m,$ where $\alpha\in(m-1,m)$ and $D_{+}^{\alpha}$ is the Rieman-Liouville derivative, is well-posed. For more information about recent results on the Cauchy problem for abstract fractional differential-operator equations, we refer the reader to [3, 7, 13]; and for a recent mathematical treatment of the distributed fractional order differential equations to papers [15, 19, 24]. These approaches to fractional evolution equations, including Bazhlekova [3], Mainardi et al. [18], and Kochubei [14], provide rigorous solution representations and fundamental solutions. However, these methods do not produce a Duhamel-type reduction: the inhomogeneous solution is not expressed as a superposition of homogeneous problems initiated by fractional impulses. The fractional analog of Duhamel’s principle, in the particular case of fractional order partial differential equations with a single ”time-fractional” term in the evolution equation, was obtained in [25, 26], for distributed order evolution equations in [27]. See also [28, 29] for the survey of the topic and [1, 2, 5, 6, 10, 12, 17, 20, 21, 22, 29, 31] for various applications of the fractional Duhamel principle. As regards evolution equations involving the fractional derivative in the sense of Riemann–Liouville with classical initial (Cauchy) conditions, such problems are ill-posed. Consequently, the notion of an “initial impulse,” discussed above, is not applicable. Nevertheless, one key feature of the Duhamel principle can be retained. Specifically, in an appropriate formulation, the same solution operator associated with the homogeneous problem can also be used to solve the nonhomogeneous problem.
@@
The operator $A\circ\mathfrak{D}^{\mathscr{A}}$ on the left-hand side of system (2) is defined as the Hadamard (entry-wise) product of the matrix $A=(a_{j,k})_{j,k=1}^{m}$ and the matrix-valued operator $\mathfrak{D}^{\mathscr{A}}=(D^{\alpha_{j,k}})_{j,k=1}^{m},$ where $D^{\alpha_{j,k}}$ denotes the Riemann–Liouville or Caputo–Djrbashian fractional derivative of order $\alpha_{j,k}.$ That is, the operator on the left hand side of (2) reads:
@@
The precise conditions on the matrix $A,$ the matrix-valued operator $F,$ and the vector-valued functions $H(t)$ and $\Phi$ that ensure the existence and uniqueness of a solution to problems (2) and (3) will be specified in Section 2, along with a review of the classical Duhamel principle and the basic reference vector spaces used in this paper. In Section 3 we present the main results: an abstract fractional analog of Duhamel’s principle for fully coupled systems together with several illustrative examples. Since the system is fully coupled, the associated fractional impulses are expected to be coupled as well, and their form of coupling is also addressed.
@@ 3. The fractional Duhamel principle for fully coupled systems
If the system coefficients satisfy $a_{ij}=0$ for $i\neq j,$ the system becomes decoupled in memory, and the history of each component evolves independently of the others. In this case, $Q(t)\equiv 0$ and the fractional impulses depend only on the history of the corresponding components. Consequently, we obtain the following theorem for fractional vector-order systems.
Theorem 12.
@@
Conclusion similar to Theorem 12 is valid in the case of systems given by the fractional derivatives in the sense of Riemann-Liouville. In the special case $m=1,$ Theorems 9 and 10 recover the fractional Duhamel principle established in [25, 26].
@@
where $H\in AC^{(}[0,T],L_{1}(\mathbb{R}))^{2}$ and $\Phi(x)\in L_{1}(\mathbb{R})\times L_{1}(\mathbb{R}).$ In equation (31) the time-fractional orders are $\alpha_{11}=1/2,\alpha_{22}=1/3$ and coefficients $a_{12}=a_{21}=0,$ implying $g_{1}(t)=g_{2}(t)=0.$ Hence, in accordance with Theorem 12 the solution of the Cauchy problem (31)-(32) has the representation